OpenStudy (anonymous):

How many grams of nitrogen are there in 5.0 grams of baking powder-(NH4)2CO3

OpenStudy (anonymous):

2 nitrogen - 14.01*2=28 grams then

OpenStudy (anonymous):

ali000000 how can there be 28 g of nitrogen if he has only 5 g sample?! now this is how i would do it: first take mass (molar mass) of the molecule and just divide mass (molar mass) of nitrogen in that molecule (so it is 2*14,01) and divide those two numbers to get percentage of N in compound and than that you multiply that percentage with mass you got (5 g). so it is: 2*M(N) / M((NH4)2CO3 = ??? and ??? * 5 g = m(N) in that given sample...

OpenStudy (anonymous):

///first take mass (molar mass) of the molecule and just divide mass (molar mass) of nitrogen in that molecule (so it is 2*14,01)/// first take mass (molar mass) of the molecule and just (multiply )mass (molar mass) of nitrogen in that molecule (so it is 2*14,01), right? correct sir @kryten